A light bulb produced by a factory is known to be 600 hours or more (undamaged) with a probability of 0.95. A sample of 10 lamps was purchased. Make a probability distribution table for the number of undamaged lights X. Calculate the expectation / mean µ and the standard deviation of X.
1
Expert's answer
2020-04-08T03:38:18-0400

Given n = 10 the number of purchased lamps and the probabilities of undamaged lamps is p, where p = 0.95. Let X be a random variable, which describes the number x of undamaged lamps purchased. The random variable X corresponds to the Binomial distribution with parameters x<=n and p (see Wikipedia for example). Thus

"Pr[X=x] = \binom{n}{x}p^x (1-p)^{n-x}" . The mean "\mu" and the standard deviation "\sigma" of X given by

"\mu = n*p = 9.5" ;

"\sigma =\sqrt{n*p*(1-p)}=\\\sqrt{10*0.95*(1-0.95)} \approx 0.689" .

The table is depicted below